A father is five years more than four times his son's age. In five years, he will be three times his son's age. How old is the son now?
Found 2 solutions by Nate, geek_angelz:
Answer by Nate(3500)

A father is five years more than four times his son's age.
f = 5 + 4s
In five years, he will be three times his son's age. How old is the son now?
f + 5 = 3(s + 5)
Plug:
5 + 4s + 5 = 3s + 15
s = 5
f = 5 + 20 = 25
s + 5 is now ... so, the son is 10 years old.

Answer by geek_angelz(1)